Nash Inc. uses a calendar year for financial reporting. The company is authorized to issue 8,810,000 shares of $12 par common stock. At no time has Nash issued any potentially dilutive securities. Listed below is a summary of Nash’s common stock activities.
1. Number of common shares issued and outstanding at December 31, 2018

1,850,000

2. Shares issued as a result of a 12% stock dividend on September 30, 2019

222,000

3. Shares issued for cash on March 31, 2020

2,180,000

Number of common shares issued and outstanding at December 31, 2020

4,252,000

4. A 2-for-1 stock split of Nash’s common stock took place on March 31, 2021
Compute the weighted-average number of common shares used in computing earnings per common share for 2019 on the 2020 comparative income statement.

shares

LINK TO TEXT

Compute the weighted-average number of common shares used in computing earnings per common share for 2020 on the 2020 comparative income statement.

shares

LINK TO TEXT

Compute the weighted-average number of common shares to be used in computing earnings per common share for 2020 on the 2021 comparative income statement.

shares

LINK TO TEXT

Compute the weighted-average number of common shares to be used in computing earnings per common share for 2021 on the 2021 comparative income statement.

shares
Answer #1

Solution A:

Weighted average number of common shares to be used in computing earnings per common share for 2019 on the 2020 comparative income statement =

Outstanding shares at beginning of 2019 restated with stock dividend regardless of where in the year the stock dividends occurs

= 1850000*112% = 2072000 shares

Solution B:

Weighted average number of common shares to be used in computing earnings per common share for 2020 on the 2020 comparative income statement:

Outstanding shares at beginning of 2020 + Weighted average nos of shares issued during the year

= 2072000 + 2180000*9/12 = 3707000 shares

Solution C:

Weighted average number of common shares to be used in computing earnings per common share for 2020 on the 2021 comparative income statement.

= 2020 weighted average nos of share outstanding * retrospective adjustment of stock split

= 3707000 * 2 = 7414000 shares

Solution D:

Weighted average number of common shares to be used in computing earnings per common share for 2021 on the 2021 comparative income statement:

Nos of outstanding shares at beginning of 2021 * retrospective adjustment of stock split

= 4252000*2 = 8504000 shares